OpenStudy (camerondoherty):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12 3y + 12 -3y + 12 3y - 12 -3y - 12

OpenStudy (anonymous):

@jim_thompson5910 he should help u :)

jimthompson5910 (jim_thompson5910):

Solve for x in -x - 3y = -12 to get -x - 3y = -12 -x - 3y+3y = -12+3y ... Add 3y to both sides. -x + 0y = -12 + 3y -x + 0 = -12 + 3y -x = -12 + 3y -x = 3y -12 x = -3y + 12 ... change the sign of every term to go from -x to x

OpenStudy (camerondoherty):

Thank You and sorry i wasnt on wifi crashed

jimthompson5910 (jim_thompson5910):

that's ok and yw

OpenStudy (anonymous):

thx for helping her @jim_thompson5910

OpenStudy (camerondoherty):

wait which one is it? those last two just confused me @jim_thompson5910

OpenStudy (camerondoherty):

brb gonna go eat btw

jimthompson5910 (jim_thompson5910):

the answer is there

jimthompson5910 (jim_thompson5910):

it's the last line

OpenStudy (camerondoherty):

oh ok thank you!

jimthompson5910 (jim_thompson5910):

you're welcome